OpenStudy (daniellelovee):

3x^(2/3)-8x^(1/3)-3=0 is the solution only 27?

OpenStudy (sparrow2):

-1/27 i think too

OpenStudy (welshfella):

let x^ (1/3) = y then x^(2/3) = y^2

OpenStudy (welshfella):

so we have the equation 3y^2 - 8y - 3 = 0 can you solve this for y?

OpenStudy (daniellelovee):

y=3, -1/3

OpenStudy (welshfella):

yes now you need to solve 3 = x^(1/3) and -1/3 = x^(1/3)

OpenStudy (welshfella):

just cube both sides of the equations

OpenStudy (daniellelovee):

x=27

OpenStudy (daniellelovee):

and x=-1/27?

OpenStudy (welshfella):

yes and ?

OpenStudy (welshfella):

yes 27 , -1/27

OpenStudy (daniellelovee):

alright thank you so much :)

OpenStudy (welshfella):

yw